# Interpretable and Steerable Concept Bottleneck Sparse Autoencoders

### v1

@@ Abstract
Sparse autoencoders (SAEs) promise a unified approach for mechanistic interpretability, concept discovery, and model steering in LLMs and LVLMs. However, realizing this potential requires that the learned features be both interpretable and steerable. To that end, we introduce two new computationally inexpensive interpretability and steerability metrics and conduct a systematic analysis on LVLMs. Our analysis uncovers two observations; (i) a majority of SAE neurons exhibit either low interpretability or low steerability or both, rendering them ineffective for downstream use; and (ii) due to the unsupervised nature of SAEs, user-desired concepts are often absent in the learned dictionary, thus limiting their practical utility. To address these limitations, we propose Concept Bottleneck Sparse Autoencoders (CB-SAE)—a novel post-hoc framework that prunes low-utility neurons and augments the latent space with a lightweight concept bottleneck aligned to a user-defined concept set. The resulting CB-SAE improves interpretability by +32.1% and steerability by +14.5% across LVLMs and image generation tasks. We will make our code and model weights available.
footnotetext: This work was performed under the auspices of the U.S. Department of Energy by the Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344, Lawrence Livermore National Security, LLC. and was supported by the LLNL-LDRD Program under Project No. 25-SI-001. LLNL-CONF-2013863
@@ 1 Introduction
Sparse autoencoders (SAEs) [karvonen2025saebench, rajamanoharan2024jumping, saelens2024] have emerged as a foundational approach for mechanistically interpreting deep neural networks. By mapping dense, polysemantic activations into sparse, overcomplete latents, SAEs produce disentangled, monosemantic features that expose internal structure and support model behavior analysis. Deployed on large language models (LLMs) [lieberum2024gemma], large-scale vision models [stevens2025saevision], and even with large vision-language models (LVLMs) [pach2025sparse] like LLaVA [zhu2024llava], SAEs promise a unified route to concept discovery and downstream model steering. However, realizing this promise requires that SAE features not only align with semantically meaningful, human-understandable concepts but also causally influence model behavior i.e., that they are both interpretable and steerable.
@@
Recent work on SAEs [arad2025saes, wang2025does] in the context of LLMs shows that interpretability does not guarantee steering effectiveness, i.e. features that activate strongly for a human-understandable concept may fail to control it when intervened upon [arad2025saes]. Although this trade-off has been observed in language models, its presence and implications in vision encoders and LVLMs remain largely unexplored. To investigate this in the LVLM setting, we conducted an empirical study of SAEs trained on activations from the CLIP [radford2021learning] vision encoder. We introduced metrics to quantify both interpretability and steerability at the neuron level and analyze their alignment across the SAE’s latent space. Our findings reveal that only about 19% of SAE neurons exhibit both high interpretability and steerability. Moreover, despite the SAE’s large dictionary size (65,536 neurons), it fails to represent 27–45% of concepts drawn from established ImageNet-derived benchmarks [subramanyam2024decider], even when trained on the corresponding data. This highlights the inability of unsupervised SAEs to cover user-specific concepts reliably.
@@
Figure 2: A. Our CB-SAE and baseline SAE can steer multiple downstream models like large vision-language models (LLaVA [liu2024improved]) or image generative models (UnCLIP [Rombach_2022_CVPR]). B. Examples of steering LLaVA and UnCLIP when using unit vector steering (zeroing out all SAE/CB-SAE neurons except the selected concept).
A counterpart to SAEs are Concept Bottleneck Models (CBMs) [koh2020concept, yuksekgonul2023posthoc, oikarinen2023labelfree, srivastava2024vlg], which approach concept learning from a supervised perspective. CBMs explicitly train a model to predict a fixed set of human-interpretable concepts by introducing a bottleneck layer that mediates the final prediction. This enables guaranteed concept coverage, but limits CBMs to predefined concepts, preventing discovery of novel features unlike SAEs. These complementary strengths highlight the need for a unified framework that combines CBMs’ controllability with SAEs’ discovery capabilities.
Motivated by these observations, we propose Concept Bottleneck Sparse Autoencoders (CB-SAE) – a unified framework that combines the unsupervised discovery capabilities of SAEs with the controllability of concept bottlenecks. We begin by pruning SAE features that lack interpretability and steerability, and then augment the resulting latent space with a lightweight CB autoencoder [kulkarni2025interpretable], trained to align with a user-specified concept set (Fig. 1B). The model is optimized using tailored loss functions that preserve reconstruction fidelity, interpretability, and steerability. As a result, our CB-SAE produces latent features that are both semantically meaningful and causally effective.
We evaluate CB-SAE on two challenging downstream tasks: controlled text generation via vision–language models (LLaVA-1.5-7B [liu2023visual], LLaVA-MORE [cocchi2025llava]) and controlled image synthesis using UnCLIP [Rombach_2022_CVPR]. CB-SAE consistently outperforms standard SAEs, with average gains of +32.1% in interpretability and +14.5% in steerability across all models and metrics. This improved performance is also shown qualitatively in Fig. 2B, where the retained SAE and CB neurons consistently outperform discarded SAE neurons w.r.t. steerability. To our knowledge, CB-SAE is the first framework to unify sparse autoencoders with concept bottleneck models, enabling robust interpretation and control of vision representations across modalities and architectures.
@@ 2 Related Work
Sparse Autoencoders. SAEs aim to discover interpretable features in neural networks by learning overcomplete decompositions of activations [makhzani2013ksae]. Recent work [bricken2023monosemanticity, huben2024sparse] showed SAEs can decompose LLM representations into monosemantic features. Various architectural innovations improved SAEs, like Batch-Top-$k$ sparsity [bussmann2024batchtopk], JumpReLU [rajamanoharan2024jumping], and Matryoshka SAEs [bussmann2025learning] with multi-level feature hierarchies. Large-scale efforts trained LLM SAEs across multiple layers and models [lieberum2024gemma, gao2025scaling], with systematic benchmarks [karvonen2025saebench]. However, we uncover two key limitations of SAEs: their unsupervised training does not guarantee the discovery of user-desired concepts, and many SAE neurons exhibit low interpretability or utility in downstream steering [arad2025saes].
Concept Bottleneck Models. CBMs [koh2020concept, yuksekgonul2023posthoc] provide a framework for building interpretable models by constraining predictions through a human-understandable concept layer, enabling both interpretation and steering. This approach has been extended to label-free settings [oikarinen2023labelfree], enhanced with vision-language guidance [yan2023learning, yang2023language, srivastava2024vlg], applied to image generative models [ismail2024concept, kulkarni2025interpretable] as well as LLMs [sun2025concept]. Our work bridges SAEs and CBMs into our novel CB-SAE, combining the expressiveness of overcomplete feature decomposition with user-specified concepts, steerability, and interpretability of concept-guided learning. A concurrent work, AlignSAE [yang2025alignsae], independently devised a similar approach to introduce supervised concepts in SAEs. They attempt to disentangle the supervised concepts from the unsupervised SAE neurons with an orthogonality loss, while our approach explicitly prunes the low utility SAE neurons and only introduces the supervised concepts absent from the retained SAE neurons. Further, AlignSAE focuses on text-based LLM SAEs while we focus on vision SAEs for multimodal LLMs and image-to-image generative models.
SAEs for Vision and Vision-Language Models. Recent work showed that SAEs can learn interpretable, monosemantic features in vision models [stevens2025saevision] as well as vision-language models [pach2025sparse, zaigrajew2025interpreting]. Another line of work [venhoff2025visual, neotowards, papadimitriou2025interpreting] investigated how visual information maps to language feature spaces via SAEs for cross-modal interpretability [nasiri2025sparc, lou2025saevlm]. However, these approaches typically neither address the challenges of ensuring discovered features are both interpretable and steerable, nor do they guarantee the discovery of user-specified concepts. Our CB-SAE addresses both limitations through post-hoc pruning and concept-bottleneck training.
@@ 3 Background
SAE preliminaries. Let $v=f_{l}(x)\in\mathbb{R}^{d}$ denote the dense activations from layer $l$ of a deep pre-trained vision model (e.g., CLIP image encoder [radford2021learning]) $f$ for an input image $x\in\mathbb{X}$. Here $d$ denotes the activation dimension and $\mathbb{X}$ corresponds to the space of images. SAEs decomposes the polysemantic activations $v$ into sparse, overcomplete latent representations $z\in\mathbb{R}^{\omega}$ ($\frac{\omega}{d}>>1$) with the aim of associating every unit in $z$ to distinct, interpretable concepts. Here $\frac{\omega}{d}$ corresponds to the expansion factor of the SAE [gao2025scaling]. Formally, an SAE is parameterized by a linear encoder $E_{\text{sae}}\in\mathbb{R}^{\omega\times d}$, a linear decoder $D_{\text{sae}}\in\mathbb{R}^{d\times\omega}$ , a shared bias term $b\in\mathbb{R}^{d}$, and a non-linear activation function $\sigma_{\text{sae}}:\mathbb{R}^{\omega}\to\mathbb{R}^{\omega}$:
@@
Measuring SAE interpretability. After training, SAEs are typically evaluated using reconstruction fidelity or sparsity [karvonen2025saebench, lou2025saevlm]. However, these metrics do not directly quantify interpretability which is namely the extent to which individual SAE neurons correspond to human-understandable concepts. While existing work [pach2025sparse, kim2025revelio] relies on manual inspection of top-$k$ activating inputs or autointerpretability scores [bills2024automated] that depend on external language models or measuring the monosemanticity [pach2025sparse], they are often subjective, computationally expensive, and difficult to scale. To this end, we leverage a popular neuron interpretability tool CLIP-Dissect [oikarinen2023clip], which utilizes a user-specified concept set $\mathcal{C}$ and a pretrained vision-language model to assign each neuron $j$ of an SAE to a human-interpretable text concept $c_{j}$. This approach used first time in the context of SAEs is computationally inexpensive and scalable. Please refer to Appendix for further details on CLIP-Dissect.
@@
Measuring SAE steerability. Beyond interpretability, SAEs have been shown to enable controllable manipulation of model behavior across language [arad2025saes], large-scale vision [stevens2025saevision], and large vision-language models [pach2025sparse, lou2025saevlm] such as LLaVA [liu2024improved]. Steerability refers to the ability to influence model outputs through targeted modifications of SAE neuron activations, thereby inducing semantically consistent changes [arad2025saes]. It is typically quantified by measuring the alignment between the steered output and the concept label of the intervened neuron. Since, in our study, the base model $f$ is a vision-only encoder, we employ a downstream generative model to evaluate the effect of SAE latent interventions.
@@
Following [pach2025sparse], we adopt LLaVA [liu2024improved], which maps an image–text pair $(x,t)$ to a text output $o$. In LLaVA, the vision encoder $f$ (e.g. CLIP [radford2021learning]) produces visual tokens $\{v_{i}\}_{i=1}^{N}$, which are projected by an adapter into the LLM’s word embedding space, where they are combined with prompt tokens to generate $o$. Similar to [pach2025sparse], to probe steerability, we use a white image with the prompt “What is shown in this image? Use exactly one word!”. For a given target neuron $j\!\in\![\omega]$ of the SAE, we overwrite its activation across all tokens (from this white image) with a fixed value $\alpha$, reconstruct the modified latents $\{\tilde{v}_{i}\}$ through the SAE decoder, and feed them into LLaVA to produce the steered output $\tilde{o}_{j}$.
@@
We then compute the cosine similarity between $\tilde{o}_{j}$ and the neuron’s CLIP-Dissect-assigned concept $c_{j}$ in a sentence-transformer embedding space [reimers-2019-sentence-bert]. Higher similarity indicates greater steerability, as the neuron reliably drives the output toward its associated concept. Unlike [pach2025sparse], which compared steered outputs to top-activating images in CLIP’s image-text space, our method compares $o_{j}$ with concepts identified by CLIP-Dissect, which produces these descriptions by aggregating activations across all images thus yielding a more robust, semantically grounded steerability metric. Note that while steering (image,text)-to-text LLaVA [liu2024improved] is one way to compute a steerability metric, a similar metric can be computed using an image-to-image generator like UnCLIP [Rombach_2022_CVPR] (see Appendix for analysis with UnCLIP).
@@ 4 Interpretability vs Steerability in SAEs
In this section, we empirically analyze SAEs from two complementary perspectives: (i) their capacity to capture interpretable concepts and (ii) their ability to steer model outputs and quantify their trade-offs. While these two measures i.e., interpretability and steerability are related, they capture distinct yet complementary aspects of SAE behavior [wang2025does]. In practice, interpretable neurons may not be steerable if their activations are weakly causal or entangled, while steerable neurons may encode abstract features misaligned with user objectives [arad2025saes, wang2025does]. The insights from this analysis on large vision-language models motivate our hybrid framework, introduced in Sec. 5, which integrates SAEs with principles from concept bottleneck models [koh2020concept, oikarinen2023labelfree, srivastava2024vlg].
@@
Setup. We train a Matryoshka Batch Top-$k$ SAE [bussmann2024batchtopk, zaigrajew2025interpreting] with $\omega$ = 65536 and an expansion factor of 64 on layer $l$ = 22 of the CLIP-ViT-L/14-336 vision encoder [radford2021learning], following the setup of [pach2025sparse], using the ImageNet-1K dataset [deng2009imagenet] (see Appendix for more details and results with other layers and models). As CLIP-Dissect requires a predefined concept set $\mathcal{C}$, we employ the Broden dataset [netdissect2017], ($|\mathcal{C}|$=1197) which provides both low-level attributes and object-level visual concepts. We then compute interpretability and steerability scores for each SAE neuron as described in Sec. 3.
Figure 4: Pipeline for building CB-SAE. Step 1. A baseline SAE is trained and Step 2. evaluated with CLIP-Dissect and downstream steering to obtain interpretability and steerability scores per SAE neuron. Step 3. The $M$ least interpretable and steerable neurons are pruned by deleting the corresponding SAE weights. Step 4. We train the CB-SAE with frozen, pruned SAE weights with three objectives: A. recover the reconstruction ability lost by pruning using $\mathcal{L}_{r}$, B. incorporate the user-specified concept set with $\mathcal{L}_{\text{int}}$, and C. promote steerability with a cyclic reconstruction loss $\mathcal{L}_{\text{st}}$.
@@
Expt. 2: Can SAEs represent all user-specified concepts? A key question is whether the SAE can represent all concepts within a given concept set, thereby supporting both human interpretability and controllable model behavior. Although the SAE contains 65,536 neurons—far exceeding the size of standard concept sets—its ability to represent concepts varies considerably with the diversity and complexity of the set. Using CLIP-Dissect, we evaluate the coverage of unique concepts across multiple concept sets:
  • Broden [netdissect2017]: 1,153/1,197 (96.3%)
  • VLG-CBM [srivastava2024vlg]: 3,445/4,729 (72.8%)
  • DECIDER [subramanyam2024decider]: 4,333/7,827 (55.3%)
  • 3k common English words [oikarinen2023labelfree]: 1,857/3,000 (61.9%)
  • 20k common English words [oikarinen2023labelfree]: 5,596/20,000 (28.0%)
@@
Our analysis reveals two key requirements: (i) expanding the SAE latent space to capture a broader range of semantically distinct concepts while remaining effective for downstream tasks such as steering, and (ii) enabling explicit user specification of concepts within the SAE. Simply pruning neurons with low interpretability and steerability degrades reconstruction fidelity. To address this and enable user-specified concepts, we propose to train a concept bottleneck autoencoder [koh2020concept, kulkarni2025interpretable] alongside the retained SAE. This hybrid framework combines (supervised) concept alignment with (unsupervised) discovery, restoring reconstruction fidelity while enhancing concept coverage and steerability.

### v2

@@ Abstract
Sparse autoencoders (SAEs) promise a unified approach for mechanistic interpretability, concept discovery, and model steering in LLMs and LVLMs. However, realizing this potential requires learned features to be both interpretable and steerable. To that end, we introduce two new computationally inexpensive interpretability and steerability metrics for a systematic analysis of LVLM SAEs. This uncovers two observations; (i) a majority of SAE neurons exhibit either low interpretability or low steerability or both, rendering them ineffective for downstream use; and (ii) user-desired concepts are often absent in the SAE, thus limiting their practical utility. To address these limitations, we propose Concept Bottleneck Sparse Autoencoders (CB-SAE) [^1] Code: github.com/Trustworthy-ML-Lab/CB-SAE—a novel post-hoc framework that prunes low-utility neurons and augments the latent space with a lightweight concept bottleneck aligned to a user-defined concept set. The resulting CB-SAE improves interpretability by +32.1% and steerability by +14.5% across LVLMs and image generation tasks.
@@ 1 Introduction
Sparse autoencoders (SAEs) [21, 43, 5] have emerged as a foundational tool for mechanistic interpretability, mapping dense polysemantic activations into sparse monosemantic latents in large language models (LLMs) [27], vision models [48, 49, 16, 53], and large vision-language models (LVLMs) [37]. However, realizing this promise requires SAE features to be semantically meaningful and causally effective, i.e. interpretable and steerable respectively.
@@
Recent work on SAEs [1, 56] in the context of LLMs shows that interpretability does not guarantee steering effectiveness, i.e. features that activate strongly for a human-understandable concept may fail to control it when intervened upon [1, 57, 21, 20]. Although this trade-off has been observed in language models, its presence and implications in vision encoders and LVLMs remain largely unexplored. To investigate this in the LVLM setting, we conducted an empirical study of SAEs trained on activations from the CLIP [42] vision encoder. We introduced metrics to quantify both interpretability and steerability at the neuron level and analyze their alignment across the SAE’s latent space. Our findings reveal that only about 19% of SAE neurons exhibit both high interpretability and steerability. Moreover, despite the SAE’s large dictionary size (65,536 neurons), it fails to represent 27–45% of concepts drawn from established ImageNet-derived benchmarks [50], even when trained on the corresponding data. This highlights the inability of unsupervised SAEs to cover user-specific concepts reliably.
@@
Figure 2: A. Our CB-SAE and baseline SAE can steer multiple downstream models like large vision-language models (LLaVA [28]) or image generative models (UnCLIP [46]). B. Examples of steering LLaVA and UnCLIP when using unit vector steering (zeroing out all SAE/CB-SAE neurons except the selected concept).
A counterpart to SAEs are Concept Bottleneck Models (CBMs) [23, 62, 34, 47, 51, 25], with concept learning from a supervised perspective. CBMs explicitly train a model to predict a fixed set of human-interpretable concepts by introducing a bottleneck layer that mediates the final prediction. This enables guaranteed concept coverage, but limits CBMs to predefined concepts, preventing discovery of novel features unlike SAEs. These complementary strengths highlight the need for a unified framework that combines CBMs’ controllability with SAEs’ discovery capabilities.
Motivated by these observations, we propose Concept Bottleneck Sparse Autoencoders (CB-SAE) – a unified framework that combines the unsupervised discovery capabilities of SAEs with the controllability of concept bottlenecks. We begin by pruning SAE features that lack interpretability and steerability, and then augment the resulting latent space with a lightweight CB autoencoder [25], trained to align with a user-specified concept set (Fig. 1B). The model is optimized using tailored loss functions that preserve reconstruction fidelity, interpretability, and steerability. As a result, our CB-SAE produces latent features that are both semantically meaningful and causally effective.
We evaluate CB-SAE on two challenging downstream tasks: controlled text generation via vision–language models (LLaVA-1.5-7B [29], LLaVA-MORE [12]) and controlled image synthesis using UnCLIP [46]. CB-SAE consistently outperforms standard SAEs, with average gains of +32.1% in interpretability and +14.5% in steerability across all models and metrics. This improved performance is also shown qualitatively in Fig. 2B, where the retained SAE and CB neurons consistently outperform discarded SAE neurons w.r.t. steerability. To our knowledge, CB-SAE is the first framework to unify sparse autoencoders with concept bottleneck models, enabling robust interpretation and control of vision representations across modalities and architectures.
@@ 2 Related Work
Sparse Autoencoders. SAEs aim to discover interpretable features in neural networks by learning overcomplete decompositions of activations [31]. Recent work [7, 18] showed SAEs can decompose LLM representations into monosemantic features. Various architectural innovations improved SAEs, like Batch-Top-$k$ sparsity [8], JumpReLU [43], and Matryoshka SAEs [9] with multi-level feature hierarchies [13]. Large-scale efforts trained LLM SAEs across multiple layers and models [27, 17], with systematic benchmarks [21]. However, we uncover two key limitations of SAEs: they do not guarantee the discovery of user-desired concepts, and many SAE neurons exhibit low interpretability or utility in downstream steering [1, 16, 57, 21, 20, 59].
Concept Bottleneck Models. CBMs [23, 62] provide a framework for building interpretable models by constraining predictions through a human-understandable concept layer, enabling both interpretation and steering. This approach has been extended to label-free settings [34], enhanced with vision-language guidance [58, 61, 47], applied to image generative models [19, 25] as well as LLMs [51]. DN-CBM [44] used an SAE encoder as a concept bottleneck layer, and trained a linear layer to classify using the SAE concepts. However, their approach is limited to image classification and susceptible to SAE limitations highlighted in our work. A concurrent work, AlignSAE [60], independently devised a similar approach to introduce supervised concepts in SAEs. They attempt to disentangle the supervised concepts from the unsupervised SAE neurons with an orthogonality loss, while our approach explicitly prunes the low utility SAE neurons and only introduces the supervised concepts absent from the retained SAE neurons. Further, AlignSAE focuses on text-based LLM SAEs while we focus on vision SAEs for multimodal LLMs and image-to-image generative models. Our work bridges SAEs and CBMs into our novel CB-SAE, combining the expressiveness of overcomplete feature decomposition with user-specified concepts, steerability, and interpretability of concept-guided learning.
SAEs for Vision and Vision-Language Models. Recent work showed that SAEs can learn interpretable, monosemantic features in vision models [49] as well as vision-language models [37, 63]. SpLiCE [3] interprets CLIP with sparse concept embeddings using supervised reconstruction and a sparsity constraint, but does not support unsupervised concept discovery unlike SAE/CB-SAE. Another line of work [55, 33, 38] investigated how visual information maps to language via SAEs for cross-modal interpretability [32, 30]. However, prior works neither address the challenges of ensuring discovered features are both interpretable and steerable, nor guarantee the discovery of user-specified concepts. Our CB-SAE addresses both limitations through post-hoc pruning and concept bottleneck training.
@@ 3 Background
SAE preliminaries. Let $v=f_{l}(x)\in\mathbb{R}^{d}$ denote the dense activations from layer $l$ of a deep pre-trained vision model (e.g., CLIP image encoder [42]) $f$ for an input image $x\in\mathbb{X}$. Here $d$ denotes the activation dimension and $\mathbb{X}$ corresponds to the space of images. SAEs decomposes the polysemantic activations $v$ into sparse, overcomplete latent representations $z\in\mathbb{R}^{\omega}$ ($\frac{\omega}{d}\gg 1$) with the aim of associating every unit in $z$ to distinct, interpretable concepts. Here $\frac{\omega}{d}$ corresponds to the expansion factor of the SAE [17]. Formally, an SAE is parameterized by a linear encoder $E_{\text{sae}}\in\mathbb{R}^{\omega\times d}$, a linear decoder $D_{\text{sae}}\in\mathbb{R}^{d\times\omega}$ , a shared bias term $b\in\mathbb{R}^{d}$, and a non-linear activation function $\sigma_{\text{sae}}:\mathbb{R}^{\omega}\to\mathbb{R}^{\omega}$:
@@
Measuring SAE interpretability. After training, SAEs are typically evaluated using reconstruction fidelity or sparsity [21, 30]. However, these metrics do not directly quantify interpretability which is namely the extent to which individual SAE neurons correspond to human-understandable concepts. While existing work [37, 22] relies on manual inspection of top-$k$ activating inputs or autointerpretability scores [4, 40] that depend on external language models or measuring the monosemanticity [37], they are often subjective, computationally expensive, and difficult to scale. To this end, we leverage a popular neuron interpretability tool CLIP-Dissect [35], which utilizes a user-specified concept set $\mathcal{C}$ and a pretrained vision-language model to assign each neuron $j$ of an SAE to a human-interpretable text concept $c_{j}$. Our interpretability score is the similarity score computed in CLIP-Dissect. This approach is computationally inexpensive, scalable, and is used for the first time with SAEs. Neuron interpretability scores have also been useful for downstream tasks like adversarial defense [24]. Please refer to Appendix for further details on CLIP-Dissect.
@@
Measuring SAE steerability. Beyond interpretability, SAEs have been shown to enable controllable manipulation of model behavior across language [1], large-scale vision [49], and large vision-language models [37, 30] such as LLaVA [28]. Steerability refers to the ability to influence model outputs through targeted modifications of SAE neuron activations, thereby inducing semantically consistent changes [1]. It is typically quantified by measuring the alignment between the steered output and the concept label of the intervened neuron. Since, in our study, the base model $f$ is a vision-only encoder, we employ a downstream generative model to evaluate the effect of SAE latent interventions.
@@
Following [37], we adopt LLaVA [28], which maps an image–text pair $(x,t)$ to a text output $o$. In LLaVA, the vision encoder $f$ (e.g. CLIP [42]) produces visual tokens $\{v_{i}\}_{i=1}^{N}$, which are projected by an adapter into the LLM’s word embedding space, where they are combined with prompt tokens to generate $o$. Similar to [37], to probe steerability, we use a white image with the prompt “What is shown in this image? Use exactly one word!”. For a given target neuron $j\!\in\![\omega]$ of the SAE, we overwrite its activation across all tokens (from this white image) with a fixed value $\alpha$, reconstruct the modified latents $\{\tilde{v}_{i}\}$ through the SAE decoder, and feed them into LLaVA to produce the steered output $\tilde{o}_{j}$.
@@
We then compute the cosine similarity between $\tilde{o}_{j}$ and the neuron’s CLIP-Dissect-assigned concept $c_{j}$ in a sentence-transformer embedding space [45]. Higher similarity indicates greater steerability, as the neuron reliably drives the output toward its associated concept. Unlike [37], which compared steered outputs to top-activating images in CLIP’s image-text space, our method compares $o_{j}$ with concepts identified by CLIP-Dissect, which produces these descriptions by aggregating activations across all images thus yielding a more robust, semantically grounded steerability metric. Note that while steering (image,text)-to-text LLaVA [28] is one way to compute a steerability metric, a similar metric can be computed using an image-to-image generator like UnCLIP [46] (see Appendix for analysis with UnCLIP).
@@ 4 Interpretability vs Steerability in SAEs
In this section, we empirically analyze SAEs from two complementary perspectives: (i) their capacity to capture interpretable concepts and (ii) their ability to steer model outputs and quantify their trade-offs. While these two measures i.e., interpretability and steerability are related, they capture distinct yet complementary aspects of SAE behavior [56]. In practice, interpretable neurons may not be steerable if their activations are weakly causal or entangled, while steerable neurons may encode abstract features misaligned with user objectives [1, 56]. The insights from this analysis on large vision-language models motivate our hybrid framework, introduced in Sec. 5, which integrates SAEs with principles from concept bottleneck models [23, 34, 47].
@@
Setup. We train a Matryoshka Batch Top-$k$ SAE [8, 63] with $\omega$ = 65536 and an expansion factor of 64 on layer $l$ = 22 of the CLIP-ViT-L/14-336 vision encoder [42], following the setup of [37], using the ImageNet-1K dataset [14] (see Appendix for more details and results with other layers and models). As CLIP-Dissect requires a predefined concept set $\mathcal{C}$, we employ the Broden dataset [2], ($|\mathcal{C}|$=1197) which provides both low-level attributes and object-level visual concepts. We then compute interpretability and steerability scores for each SAE neuron as described in Sec. 3.
Figure 4: Pipeline for building CB-SAE. Step 1. A baseline SAE is trained and Step 2. evaluated with CLIP-Dissect and downstream steering to obtain interpretability and steerability scores per SAE neuron. Step 3. The $M$ least interpretable and steerable neurons are pruned by deleting the corresponding SAE weights. Step 4. We train the CB-SAE with frozen, pruned SAE weights with three objectives: A. recover the reconstruction ability lost by pruning using $\mathcal{L}_{r}$, B. incorporate the user-specified concept set with $\mathcal{L}_{\text{int}}$, and C. promote steerability with a cyclic reconstruction loss $\mathcal{L}_{\text{st}}$. Outcome. CB-SAE uses both retained SAE and CB-AE for inference.
@@
Expt. 2: Can SAEs represent all user-specified concepts? A key question is whether the SAE can represent all concepts from a given concept set, thereby supporting both human interpretability and controllable model behavior. Although the SAE contains 65,536 neurons, far exceeding the size of standard concept sets, its ability to represent concepts varies considerably with the diversity and complexity of the set.
Observations. Using CLIP-Dissect, we evaluate the coverage of unique concepts across multiple concept sets:
  • Broden [2]: 1,153/1,197 (96.3%)
  • VLG-CBM [47]: 3,445/4,729 (72.8%)
  • DECIDER [50]: 4,333/7,827 (55.3%)
  • 3k common English words [34]: 1,857/3,000 (61.9%)
  • 20k common English words [34]: 5,596/20,000 (28.0%)
@@
Our analysis reveals two key requirements: (i) expanding the SAE latent space to capture a broader range of semantically distinct concepts while remaining effective for downstream tasks such as steering, and (ii) enabling explicit user specification of concepts within the SAE. Simply pruning neurons with low interpretability and steerability degrades reconstruction fidelity. To address this and enable user-specified concepts, we propose to train a concept bottleneck autoencoder [23, 25] alongside the retained SAE. This hybrid framework combines (supervised) concept alignment with (unsupervised) discovery, restoring reconstruction fidelity while enhancing concept coverage and steerability.